Line P passes through point (5,1) and has a slope of -4/5. What is the equation of line p in standard form?
Answer by mananth(16946)

Line P passes through point (5,1) and has a slope of -4/5. What is the equation of line p in standard form?
...
y=mx+b
plug the above values
1=(-4/5)*5+b
1=-4+b
1+4=b
b=5
....
y= mx+b
y=(-4/5)x+5
multiply by 5
5y=-4x+25
4x+5y-25=0
...
m.ananth@hotmail.ca